OpenStudy (anonymous):

Help please!! (Metal and fan) Using point-slope form, write the equation of the line that passes through the point (-1/2, 1/2) and has a slope of -1

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@superdavesuper

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@abb0t

OpenStudy (superdavesuper):

point and slope form is (y-y1)=m(x-x1) + c, where m is slope n c is a constant. so u got x1, y1 and m already; just need to solve for c

OpenStudy (anonymous):

Y-1/2=-x+1/2 ??

OpenStudy (anonymous):

y+1/2=-x-1/2

OpenStudy (superdavesuper):

Wait wait im sorry....i think i made a mistake.....

OpenStudy (anonymous):

or y+1/2=-x+1/2

OpenStudy (superdavesuper):

point and slope form is (y-y1)=m(x-x1) so u got it already :)

OpenStudy (superdavesuper):

sorry about my mistake @RobertoTB

OpenStudy (anonymous):

so which one is it

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

(y-y1)=m(x-x1) and just put in x1, y1 and m

OpenStudy (anonymous):

y-1/2=-x-1/2

OpenStudy (anonymous):

this one @superdavesuper

OpenStudy (superdavesuper):

yup good job! :)